Healthlabs Ltd. purchased lab equipment for $18,300 on January 1, 2020. At that time, management determined that the equipment would have a useful life of three years with no residual value and that the straight-line depreciation method would be used. On January 1, 2021, management revised its initial estimate and determined that the company would be able to use the lab equipment for a total of five years.

Determine the depreciation expense for 2021.

Depreciation expense for 2021 $Enter the depreciation expense for 2021 in dollars.
Answer #1
Book value on 01/jan 2021 :- Original Cost- Depreciation
=18300-6100=$12200
*Depreciation Expense= 18300/3=6100
Revised Useful Life : total Five Year
Remaining useful life=4 year
Depreciation expense for 2021 = Book Value / Remaining Useful life
(12200/4)=$3050

Answer will be $3050